OpenStudy (anonymous):

To prove that 1 + 2 + 3 + ... + n = n(n + 1)/1 by mathematical induction, the first step would be

OpenStudy (anonymous):

you just have to identify the first step here are the choices A. Show that 1 = 1(1 + 1)/2 B. Show that 1 + 2 = 2(2 + 1)/2 C. Show that 1 + 2 + 3 = 3(3 + 1)/2 D. Assume that 1 + 2 + ... + k = k(k + 1)/2

OpenStudy (anonymous):

pretty lost I'll just go watch a video or something..

OpenStudy (anonymous):

thanks for your help though